OpenStudy (anonymous):

Last month, Marcus borrowed $87 from his roommate. So far, he has paid him back $27. Which equation could be used to find x, the amount of money he still owes his roommate?

OpenStudy (anonymous):

87-27=x

OpenStudy (anonymous):

a$87 + x = $27 b$27x = $87 cx + $27 = $87 dx - $27 = $87

OpenStudy (anonymous):

X+27=87